OpenStudy (anonymous):

If y varies inversely as x, and y = 3 when x = 6, find the value of x when y = 18. Answer 36 6 9 1

OpenStudy (yttrium):

Just be reminded that if this is the case. y = k/x Solve for k. Then solve for x when y is 18.

OpenStudy (anonymous):

since its inverse , x1y1 = x2y2 6*3 = x2*18 ,solve the equation